# Compact Minimal Submanifolds of the Riemannian Symmetric Spaces

### v1

@@ Abstract.
In this work we construct new multi-dimensional families of compact minimal submanifolds, of the classical Riemannian symmetric spaces $\text{\bf SU}(n)/\mathbf{SO}(n)$, $\text{\bf Sp}(n)/\text{\bf U}(n)$, $\mathbf{SO}(2n)/\text{\bf U}(n)$ and $\text{\bf SU}(2n)/\text{\bf Sp}(n)$, of codimension two.
@@ 1. Introduction
This was later generalised to the study of minimal surfaces in much more general ambient manifolds via conformal harmonic immersions. The next result follows from the seminal paper [6] of Eells and Sampson from 1964. For this see also Proposition 3.5.1 of [2].
@@
In their work [1] from 1981, Baird and Eells have shown that complex-valued harmonic morphisms from Riemannian manifolds are useful tools for the study of minimal submanifolds of codimension two. This can be seen as dual to the above mentioned generalisation of the Weierstrass-Enneper representation. Harmonic morphisms are the much studied horizontally conformal harmonic maps. For an introduction to the theory of harmonic morphisms we recommend the book [2], by Baird and Wood, and the regularly updated online bibliography [11].
@@
The recent work [13], introduces a method for constructing minimal submanifolds of Riemannian manifolds, via submersions, see Theorem 3.1. Then this scheme is employed to provide compact examples in several important cases.
The main ingredients for this new procedure are the so called complex-valued eigenfunctions on the Riemannian ambient space. These are functions which are eigen both with respect to the classical Laplace-Beltrami and the so called conformality operator, see Section 2.
In the current study we continue the investigation and apply the new above mentioned method to the classical Riemannian symmetric spaces
$\text{\bf SU}(n)/\mathbf{SO}(n),\ \text{\bf Sp}(n)/\text{\bf U}(n),\ \mathbf{SO}(2n)/\text{\bf U}(n),\ \text{\bf SU}(2n)/\text{\bf Sp}(n).$
In these cases we find multi-dimensional families of compact minimal submanifolds of codimension two, see Theorems 8.2, 9.2, 10.2 and 11.2. Readers interested in further details are referred to [8].
@@ 2. Eigenfunctions and Eigenfamilies
Let $(M,g)$ be an $m$-dimensional Riemannian manifold and $T^{{\mathbb{C}}}M$ be the complexification of the tangent bundle $TM$ of $M$. We extend the metric $g$ to a complex bilinear form on $T^{{\mathbb{C}}}M$. Then the gradient $\nabla\phi$ of a complex-valued function $\phi:(M,g)\to{\mathbb{C}}$ is a section of $T^{{\mathbb{C}}}M$. In this situation, the well-known complex linear Laplace-Beltrami operator (alt. tension field) $\tau$ on $(M,g)$ acts locally on $\phi$ as follows
@@
$U/K$ $\lambda$ $\mu$ Eigenfunctions $\mathbf{SO}(n)$ $-\,\frac{(n-1)}{2}$ $-\,\frac{1}{2}$ see [14] $\text{\bf SU}(n)$ $-\,\frac{n^{2}-1}{n}$ $-\,\frac{n-1}{n}$ see [16] $\text{\bf Sp}(n)$ $-\,\frac{2n+1}{2}$ $-\,\frac{1}{2}$ see [12] $\text{\bf SU}(n)/\mathbf{SO}(n)$ $-\,\frac{2(n^{2}+n-2)}{n}$ $-\,\frac{4(n-1)}{n}$ see [15] $\text{\bf Sp}(n)/\text{\bf U}(n)$  $-\,2(n+1)$  $-\,2$ see [15] $\mathbf{SO}(2n)/\text{\bf U}(n)$  $-\,2(n-1)$ $-1$ see [15] $\text{\bf SU}(2n)/\text{\bf Sp}(n)$ $-\,\frac{2(2n^{2}-n-1)}{n}$ $-\,\frac{2(n-1)}{n}$ see [15] $\mathbf{SO}(m+n)/\mathbf{SO}(m)\times\mathbf{SO}(n)$ $-(m+n)$ $-2$ see [9] $\text{\bf U}(m+n)/\text{\bf U}(m)\times\text{\bf U}(n)$ $-2(m+n)$ $-2$ see [10] $\text{\bf Sp}(m+n)/\text{\bf Sp}(m)\times\text{\bf Sp}(n)$ $-2(m+n)$ $-1$ see [10]
@@
The next result shows that a given eigenfamily $\mathcal{E}$ on $(M,g)$ induces a large collection $\mathcal{P}_{d}(\mathcal{E})$ of such objects. This is particularly interesting in the light of Table 1.
Theorem 2.4.
[10] Let $(M,g)$ be a Riemannian manifold and the set of complex-valued functions
=$\mathcal{E}=\{\phi_{i}:M\to{\mathbb{C}}\,|\,i=1,2,\dots,n\}$
be a finite eigenfamily i.e. there exist complex numbers $\lambda,\mu\in{\mathbb{C}}$ such that for all $\phi,\psi\in\mathcal{E}$
Then the set of complex homogeneous polynomials of degree $d$
$\mathcal{P}_{d}(\mathcal{E})=\{P:M\to{\mathbb{C}}\,|\,P\in{\mathbb{C}}[\phi_{1},\phi_{2},\dots,\phi_{n}],\,P(\alpha\cdot\phi)=\alpha^{d}\cdot P(\phi),\,\alpha\in{\mathbb{C}}\}$
is an eigenfamily on $M$ such that for all $P,Q\in\mathcal{P}_{d}(\mathcal{E})$ we have
$\tau(P)=(d\,\lambda+d(d-1)\,\mu)\cdot P\ \ \text{and}\ \ \kappa(P,Q)=d^{2}\mu\cdot P\,Q.$
@@
Let $\pi:(\hat{M},\hat{g})\to(M,g)$ be a Riemannian submersion between Riemannian manifolds. Further let $\phi:(M,g)\to{\mathbb{C}}$ be a smooth function and $\hat{\phi}:(\hat{M},\hat{g})\to{\mathbb{C}}$ be the composition $\hat{\phi}=\phi\circ\pi$. Then the corresponding tension fields $\tau$ and conformality operators $\kappa$ satisfy
@@
The arguments needed here can be found in [15]
@@
Let $\pi:(\hat{M},\hat{g})\to(M,g)$ be a Riemannian submersion. For a complex-valued function smooth function $\phi:(M,g)\to{\mathbb{C}}$ let $\hat{\phi}:(\hat{M},\hat{g})\to{\mathbb{C}}$ be the composition $\hat{\phi}=\phi\circ\pi$. Then the following statements are equivalent
@@ 4. The General Linear Group $\mathbf{GL}_{n}({\mathbb{C}})$
The group of linear automorphism of ${\mathbb{C}}^{n}$ is the complex general linear group $\mathbf{GL}_{n}({\mathbb{C}})=\{z\in{\mathbb{C}}^{n\times n}\,|\,\det z\neq 0\}$ of invertible $n\times n$ matrices with its standard representation
@@ 8. The Symmetric Space $\text{\bf SU}(n)/\mathbf{SO}(n)$
The purpose of this section is to prove Theorem 8.2 and thereby construct a new multi-dimensional family of compact minimal submanifolds of the homogeneous quotient manifold $\text{\bf SU}(n)/\mathbf{SO}(n)$, which carries the structure of a compact Riemannian symmetric space. It is well-known that the natural projecton $\pi:\text{\bf SU}(n)\to\text{\bf SU}(n)/\mathbf{SO}(n)$ from the special unitary group $\text{\bf SU}(n)$ is a Riemannian submersion. This means that we can apply Corollary 2.6 in this situation.
@@
With our next result we construct a complex $(n-1)$-dimensional family
=$\mathcal{F}=\{\phi_{a}^{-1}(\{0\})\ |\ [a]\in{\mathbb{C}}P^{n-1}\}$
Theorem 8.2.
$\hat{\phi}_{a}(z)=\operatorname{trace}(Azz^{t}).$
Then $\phi_{a}$ is a submersion and the compact fibre $\phi_{a}^{-1}(\{0\})$ over $0\in{\mathbb{C}}$ is a minimal submanifold of $\text{\bf SU}(n)/\mathbf{SO}(n)$ of codimension two.
@@
For the Lie algebra $\mathfrak{su}(n)$ of $\text{\bf SU}(n)$ we have the orthogonal decomposition $\mathfrak{su}(n)=\mathfrak{so}(n)\oplus i\mathfrak{p}$ where the subspace
@@ 9. The Symmetric Space $\text{\bf Sp}(n)/\text{\bf U}(n)$
In this section we prove Theorem 9.2 and thereby construct a new multi-dimensional family of compact minimal submanifolds of the Riemannian symmetric space $\text{\bf Sp}(n)/\text{\bf U}(n)$. It is well-known that the natural projecton $\pi:\text{\bf Sp}(n)\to\text{\bf Sp}(n)/\text{\bf U}(n)$ from the quaternionic unitary group $\text{\bf Sp}(n)$ is a Riemannian submersion. This means that we can employ Corollary 2.6 in this situation.
@@
With the following result we construct a complex $(2n-1)$-dimensional family
=$\mathcal{F}=\{\phi_{a}^{-1}(\{0\})\ |\ [a]\in{\mathbb{C}}P^{2n-1}\}$
of minimal submanifolds of $\text{\bf Sp}(n)/\text{\bf U}(n)$ of codimension two. This is our main result of this section.
Theorem 9.2.
$\hat{\phi}_{a}(q)=\operatorname{trace}(Aqq^{t}).$
Then $\phi_{a}$ is a submersion and the compact fibre $\phi_{a}^{-1}(\{0\})$ over $0\in{\mathbb{C}}$ is a minimal submanifold of $\text{\bf Sp}(n)/\text{\bf U}(n)$ of codimension two.
@@
Let us assume that $q\in\text{\bf Sp}(n)$ is a critical point for the function $\hat{\phi}_{a}$ i.e. that the gradient
@@
Equation 9.1 implies that, for all $1\leq t\leq n$, either
@@
By plugging this into the equation 9.1 we then yield
@@ 10. The Symmetric Space $\mathbf{SO}(2n)/\text{\bf U}(n)$
In this section we prove Theorem 10.2 and thereby construct a new multi-dimensional family of compact minimal submanifolds of the Riemannian symmetric space $\mathbf{SO}(2n)/\text{\bf U}(n)$. It is well-known that the natural projecton $\pi:\mathbf{SO}(2n)\to\mathbf{SO}(2n)/\text{\bf U}(n)$ from the special orthogonal group $\mathbf{SO}(2n)$ is a Riemannian submersion. This means that we can apply Corollary 2.6 in this situation.
@@
We now state Theorem 10.2, which is our main result of this section.
Theorem 10.2.
$A=\sum_{j,\alpha=1}^{2n}a_{j}b_{\alpha}Y_{j\alpha}.$
$\hat{\phi}_{a,b}(x)=\operatorname{trace}(AxJ_{n}x^{t}).$
@@
From the definition $\hat{\phi}_{a,b}(x)=\operatorname{trace}(AxJ_{n}x^{t})$ of $\hat{\phi}_{a,b}$, it is easily seen that
@@
$(u,x_{r})(b,x_{n+s})+(u,x_{n+r})(b,x_{s})=(u,x_{s})(b,x_{n+r})+(u,x_{n+s})(b,x_{r}),$
$(u,x_{r})(b,x_{s})+(u,x_{n+s})(b,x_{n+r})=(u,x_{s})(b,x_{r})+(u,x_{n+r})(b,x_{n+s}),$
respectively. The non vanishing vector $a=u+i\,v\in{\mathbb{C}}^{2n}$ is isotropic i.e. $(a,a)=0$, hence $|u|=|v|\neq 0$. This implies that there exists an $r\in\{1,2,\dots,2n\}$ such that $(u,x_{r})\neq 0$. Without loss of generality we can assume that $(u,x_{1})\neq 0$. Then with $s=2,3,\dots n$ we yield the following system of $22n=2\,(n-1)+$ 2 equations
@@
where $I_{2n}$ is the identity matrix and $S_{n}\in{\mathbb{R}}^{2n\times 2n}$ is skew-symmetric, see Appendix A. This means that the matrix $M_{n}$ is invertible. A lengthy and rather tedious calculation shows that $M_{n}\in{\mathbb{R}}^{2n\times 2n}$ satisfies
$|\det M_{n}|=|\det\tilde{M}_{n}|=(\,(u,x_{1})^{2}+(u,x_{n+1})^{2}\,)^{n-2}\cdot|u|^{4}\neq 0.$
This implies that the non-zero vector $b\in{\mathbb{R}}^{2n}$ vanishes. This contradicts the fact that an element $x\in\hat{\phi}_{a,b}^{-1}(\{0\})$ is a critical point of the $\text{\bf U}(n)$-invariant function $\hat{\phi}_{a,b}:\mathbf{SO}(2n)\to{\mathbb{C}}$. It then follows that the compact fibre $\phi_{a,b}^{-1}(\{0\})$ is regular. According to Theorem 3.1 it is minimal as well. ∎

### v2

@@ Abstract.
In this work we construct new multi-dimensional families of compact minimal submanifolds of the classical Riemannian symmetric spaces $\text{\bf SU}(n)/\mathbf{SO}(n)$, $\text{\bf Sp}(n)/\text{\bf U}(n)$, $\mathbf{SO}(2n)/\text{\bf U}(n)$ and $\text{\bf SU}(2n)/\text{\bf Sp}(n)$ of codimension two.
@@ 1. Introduction
This was later generalised to the study of minimal surfaces in much more general ambient manifolds via harmonic conformal immersions. The next result follows from the seminal paper [6] of Eells and Sampson from 1964. For this see also Proposition 3.5.1 of [2].
@@
In their work [1] from 1981, Baird and Eells have shown that complex-valued harmonic morphisms from Riemannian manifolds are useful tools for the study of minimal submanifolds of codimension two.
@@
This can be seen as dual to the above mentioned generalisation of the Weierstrass-Enneper representation. Harmonic morphisms are the much studied horizontally conformal harmonic maps. For an introduction to the theory we recommend the book [2], by Baird and Wood, and the regularly updated online bibliography [11].
2. The Main Results
The recent work [12] introduces a method for constructing minimal submanifolds of Riemannian manifolds via submersions, see Theorem 4.1. Then this scheme is employed to provide compact examples in several important cases. The main ingredients for this new procedure are the so called complex-valued eigenfunctions on the Riemannian ambient space. These are functions which are eigen both with respect to the classical Laplace-Beltrami and the so called conformality operator, see Section 3.
In the current study we continue the investigation and apply the above mentioned method to the classical Riemannian symmetric spaces
$\text{\bf SU}(n)/\mathbf{SO}(n),\ \text{\bf Sp}(n)/\text{\bf U}(n),\ \mathbf{SO}(2n)/\text{\bf U}(n)\ \ \text{and}\ \ \text{\bf SU}(2n)/\text{\bf Sp}(n).$
In the first case we construct a complex $(n-1)$-dimensional family
$\mathcal{F}_{1}=\{\phi_{a}^{-1}(\{0\})\ |\ [a]\in{\mathbb{C}}P^{n-1}\}$
Theorem 2.1.
$\hat{\phi}_{a}(z)=\operatorname{trace}(Azz^{t}).$
Then the origin $0\in{\mathbb{C}}$ is a regular value of $\phi_{a}$, so the compact fibre $\phi_{a}^{-1}(\{0\})$ is a minimal submanifold of $\text{\bf SU}(n)/\mathbf{SO}(n)$ of codimension two.
For the Riemannian symmetric space $\text{\bf Sp}(n)/\text{\bf U}(n)$ we construct a complex $(2n-1)$-dimensional family
$\mathcal{F}_{2}=\{\phi_{a}^{-1}(\{0\})\ |\ [a]\in{\mathbb{C}}P^{2n-1}\}$
of minimal submanifolds of codimension two.
Theorem 2.2.
$\hat{\phi}_{a}(q)=\operatorname{trace}(Aqq^{t}).$
Then the origin $0\in{\mathbb{C}}$ is a regular value of $\phi_{a}$, so the compact fibre $\phi_{a}^{-1}(\{0\})$ is a minimal submanifold of $\text{\bf Sp}(n)/\text{\bf U}(n)$ of codimension two.
In the third case we construct a real $(6n-3)$-dimensional family
$\mathcal{F}_{3}=\{\phi_{a,b}^{-1}(\{0\})\ |\ [a]\in{\mathbb{C}}P^{2n-1},\ [b]\in{\mathbb{R}}P^{2n-1},\ (a,a)=(a,b)=0\}$
of minimal submanifolds of $\mathbf{SO}(2n)/\text{\bf U}(n)$ of codimension two.
Theorem 2.3.
$A=\sum_{j,\alpha=1}^{2n}a_{j}b_{\alpha}Y_{j\alpha}.$
$\hat{\phi}_{a,b}(x)=\operatorname{trace}(AxJ_{n}x^{t}).$
For the Riemannian symmetric space $\text{\bf SU}(2n)/\text{\bf Sp}(n)$ we construct a complex $(4n-2)$-dimensional family
$\mathcal{F}_{4}=\{\phi_{a,b}^{-1}(\{0\})\ |\ [a],[b]\in{\mathbb{C}}P^{2n-1},[a]\neq[b]\}$
of minimal submanifolds of codimension two.
Theorem 2.4.
$A=\sum_{j,\alpha=1}^{2n}a_{j}b_{\alpha}Y_{j\alpha}.$
$\hat{\phi}_{a,b}(z)=\operatorname{trace}(AzJ_{n}z^{t}).$
The proofs of these four results are provided below. The readers interested in further details are referred to [8].
@@ 3. Eigenfunctions and Eigenfamilies
Let $(M,g)$ be an $m$-dimensional Riemannian manifold and $T^{{\mathbb{C}}}M$ be the complexification of the tangent bundle $TM$ of $M$. We extend the metric $g$ to a complex bilinear form on $T^{{\mathbb{C}}}M$. Then the gradient $\nabla\phi$ of a complex-valued function $\phi:(M,g)\to{\mathbb{C}}$ is a section of $T^{{\mathbb{C}}}M$. In this situation, the well-known complex linear Laplace-Beltrami operator (alt. tension field) $\tau$ on $(M,g)$. In local coordinates this satisfies
@@
$G/K$ $\lambda$ $\mu$ Eigenfunctions $\mathbf{SO}(n)$ $-\,\frac{(n-1)}{2}$ $-\,\frac{1}{2}$ see [13] $\text{\bf SU}(n)$ $-\,\frac{n^{2}-1}{n}$ $-\,\frac{n-1}{n}$ see [15] $\text{\bf Sp}(n)$ $-\,\frac{2n+1}{2}$ $-\,\frac{1}{2}$ see [13] $\text{\bf SU}(n)/\mathbf{SO}(n)$ $-\,\frac{2(n^{2}+n-2)}{n}$ $-\,\frac{4(n-1)}{n}$ see [14] $\text{\bf Sp}(n)/\text{\bf U}(n)$  $-\,2(n+1)$  $-\,2$ see [14] $\mathbf{SO}(2n)/\text{\bf U}(n)$  $-\,2(n-1)$ $-1$ see [14] $\text{\bf SU}(2n)/\text{\bf Sp}(n)$ $-\,\frac{2(2n^{2}-n-1)}{n}$ $-\,\frac{2(n-1)}{n}$ see [14] $\mathbf{SO}(m+n)/\mathbf{SO}(m)\times\mathbf{SO}(n)$ $-(m+n)$ $-2$ see [9] $\text{\bf U}(m+n)/\text{\bf U}(m)\times\text{\bf U}(n)$ $-2(m+n)$ $-2$ see [10] $\text{\bf Sp}(m+n)/\text{\bf Sp}(m)\times\text{\bf Sp}(n)$ $-2(m+n)$ $-1$ see [10]
@@
Let $\pi:(\hat{M},\hat{g})\to(M,g)$ be a harmonic Riemannian submersion between Riemannian manifolds. Further let $\phi:(M,g)\to{\mathbb{C}}$ be a smooth function and $\hat{\phi}:(\hat{M},\hat{g})\to{\mathbb{C}}$ be the composition $\hat{\phi}=\phi\circ\pi$. Then the corresponding tension fields $\tau$ and conformality operators $\kappa$ satisfy
@@
The arguments needed here can be found in [14]. ∎
@@
Let $\pi:(\hat{M},\hat{g})\to(M,g)$ be a harmonic Riemannian submersion. For a complex-valued function smooth function $\phi:(M,g)\to{\mathbb{C}}$ let $\hat{\phi}:(\hat{M},\hat{g})\to{\mathbb{C}}$ be the composition $\hat{\phi}=\phi\circ\pi$. Then the following statements are equivalent
@@ 5. The General Linear Group $\mathbf{GL}_{n}({\mathbb{C}})$
The group of linear automorphisms of ${\mathbb{C}}^{n}$ is the complex general linear group $\mathbf{GL}_{n}({\mathbb{C}})=\{z\in{\mathbb{C}}^{n\times n}\,|\,\det z\neq 0\}$ of invertible $n\times n$ matrices with its standard representation
@@ 9. The Symmetric Space $\text{\bf SU}(n)/\mathbf{SO}(n)$
The purpose of this section is to prove Theorem 2.1 and thereby construct a new multi-dimensional family of compact minimal submanifolds of the homogeneous quotient manifold $\text{\bf SU}(n)/\mathbf{SO}(n)$, which carries the structure of a compact Riemannian symmetric space. It is well-known that the natural projection $\pi:\text{\bf SU}(n)\to\text{\bf SU}(n)/\mathbf{SO}(n)$ from the special unitary group $\text{\bf SU}(n)$ is a Riemannian submersion. This means that we can apply Corollary 3.5 in this situation.
@@
(Theorem 2.1) For the Lie algebra $\mathfrak{su}(n)$ of $\text{\bf SU}(n)$ we have the orthogonal decomposition $\mathfrak{su}(n)=\mathfrak{so}(n)\oplus i\mathfrak{p}$ where the subspace
@@ 10. The Symmetric Space $\text{\bf Sp}(n)/\text{\bf U}(n)$
In this section we prove Theorem 2.2, which yields a new multi-dimensional family of compact minimal submanifolds of the Riemannian symmetric space $\text{\bf Sp}(n)/\text{\bf U}(n)$. We first recall that the natural projection $\pi:\text{\bf Sp}(n)\to\text{\bf Sp}(n)/\text{\bf U}(n)$ from the quaternionic unitary group $\text{\bf Sp}(n)$ is a Riemannian submersion and thus we may now employ Corollary 3.5.
@@
(Theorem 2.2) Let us assume that $q\in\text{\bf Sp}(n)$ is a critical point for the function $\hat{\phi}_{a}$ i.e. that the gradient
@@
Equation (10.1) implies that for all $1\leq t\leq n$ either
@@
By plugging this into the equation (10.1) we then yield
@@ 11. The Symmetric Space $\mathbf{SO}(2n)/\text{\bf U}(n)$
The goal of this section is to prove Theorem 2.3, which immediately provides us with a new multi-dimensional family of compact minimal submanifolds of the Riemannian symmetric space $\mathbf{SO}(2n)/\text{\bf U}(n)$. As the natural projection $\pi:\mathbf{SO}(2n)\to\mathbf{SO}(2n)/\text{\bf U}(n)$ from the special orthogonal group $\mathbf{SO}(2n)$ is a Riemannian submersion, we can now apply Corollary 3.5.
@@
(Theorem 2.3) From the definition
$\hat{\phi}_{a,b}(x)=\operatorname{trace}(AxJ_{n}x^{t})$
of $\hat{\phi}_{a,b}$ it is easily seen that
@@
$0=(u,x_{r})(b,x_{n+s})+(u,x_{n+r})(b,x_{s})-(u,x_{s})(b,x_{n+r})-(u,x_{n+s})(b,x_{r}),$
$0=(u,x_{r})(b,x_{s})+(u,x_{n+s})(b,x_{n+r})-(u,x_{s})(b,x_{r})-(u,x_{n+r})(b,x_{n+s}),$
respectively. The non-vanishing vector $a=u+i\,v\in{\mathbb{C}}^{2n}$ is isotropic i.e. $(a,a)=0$, hence $|u|=|v|\neq 0$. This implies that there exists an $r\in\{1,2,\dots,2n\}$ such that $(u,x_{r})\neq 0$. Without loss of generality we can assume that $(u,x_{1})\neq 0$. Then with $s=2,3,\dots n$ we yield the following system of $22n=2\,(n-1)+$ 2 equations
@@
where $I_{2n}$ is the identity matrix and $S_{n}\in{\mathbb{R}}^{2n\times 2n}$ is skew-symmetric, see Appendix A. This means that the matrix $M_{n}$ is invertible. This implies that the non-zero vector $b\in{\mathbb{R}}^{2n}$ vanishes. This contradicts the fact that an element $x\in\hat{\phi}_{a,b}^{-1}(\{0\})$ is a critical point of the $\text{\bf U}(n)$-invariant function $\hat{\phi}_{a,b}:\mathbf{SO}(2n)\to{\mathbb{C}}$. It then follows that the compact fibre $\phi_{a,b}^{-1}(\{0\})$ is regular. According to Theorem 4.1 it is minimal as well. ∎